The table below provides data taken from a study that attempts to determine whether the use of electronic fetal monitoring (EFM) during labor affects the frequency of cesarean section deliveries. Of the 5824 infants included in the study, 2850 were electronically monitored and 2974 were not. The outcomes are listed in Table E1.6.

(a) Compute and compare the proportions of cesarean delivery for the two exposure groups. What would be your conclusion?

(b) Calculate the odds ratio associated with EFM exposure. Does this result support your conclusion in part (a)?

TABLE E1.6 EFM exposur No 229 Cesarean delivery Yes Total Yes No Total 358 2492 2850 587 5237 29745824 2745
Answer #1

Ans:

a)proportions of cesarean delivery for group with EFM exposure=358/2850=0.1256

proportions of cesarean delivery for group without EFM exposure=229/2974=0.0770

Proportion of Cesarean delivery for group without EFM exposure is lower than proportion of Cesarean delivery for group with EFM exposure.

b)

Odds(Cesarean delivery)=358/229=1.5633

Odds(not Cesarean delivery)=2492/2745=0.9078

Odds ratio=1.5633/0.9078=1.722

Yes,as rate of Cesarean delivery in more in the group with EFM exposure than without EFM exposure,which is indicated by OR>1